Find the sum of the infinite geometric series 36+12+4...
Answer by amarjeeth123(569)

The sum is equal to a/(1-r)
Here a=36=first term
r=Common ratio=12/36=1/3
Sum=36/(1-1/3)=36/(2/3)=36*(3/2)=54.